# Towards Full-scene Domain Generalization in Multi-agent Collaborative Bird’s Eye View Segmentation for Connected and Autonomous Driving

### v2

@@ IV-A Datasets and Metrics
To assess the domain generalization performance of our proposed method, it is essential to evaluate it across various datasets exhibiting domain discrepancies. However, in the domain of collaborative BEV segmentation, there is no public available datasets suitable for this purpose. Consequently, we have taken the initiative to synthesize four distinct datasets, each characterized with unique domain discrepancies, drawing from the OPV2V dataset xuOPV2VOpenBenchmark2022 as the foundation. These datasets include variations such as foggy-weather, rainy-weather, night conditions, and bright sunny scenarios. The specific attributes and details of these synthesized datasets are elaborated upon below.
OPV2V: In our study, we leverage the OPV2V dataset xuOPV2VOpenBenchmark2022 , a comprehensive dataset curated for integrated perception using V2V communications. This dataset, amassed through the CARLA simulator dosovitskiyCARLAOpenUrban2017 and OpenCDA xuOpenCDAOpenCooperative2021 , comprises 11464 frames of LiDAR point clouds and images, with each frame featuring the minimum of 2 and the maximum of 7 interconnected vehicles. It encompasses 73 distinct scenarios, each averaging a duration of 25 seconds.
Sunny: The initial task in our setting is driving on a sunny day. As the OPV2V dataset is collected at the daytime, so we sample a portion of the OPV2V as our sunny dataset.
Fog: The second task is driving on a foggy day. We synthesize a foggy-weather dataset with different densities of fog (e.g., dense fog and light fog) as shown in Fig. 4 and 5. The foggy weather is generated from atmospheric scattering model narasimhanVisionAtmosphere2008 .
Rain: In addition, we synthesize a rainy-weather dataset with different densities of rain (e.g., heavy rain and light rain) as shown in Fig. 6. Specifically, the dataset is generated by utilizing RainMix, which is a novel data augmentation method. In order to generate the rainy dataset, we first sample a rain map from a public dataset of real rain streaks gargPhotorealisticRenderingRain2006 . We then apply random transformations to the rain map using the RainMix technique. These transformations include rotation, zooming, translation, and shearing, which are randomly sampled and combined. Lastly, the rain maps after transformation are merged with the original source domain images.
Night: Finally, as shown in Fig. 7, we synthesize a night dataset by leveraging a day-to-night generative adversarial network (GAN) zhuUnpairedImagetoImageTranslation2020 ; arrudaCrossDomainCarDetection2019 , which can render the daytime images to night images.
@@
Domain Type
Sunny Fog
AP@IoU(%)
Vehicle Road Lane Average Vehicle Road Lane Average
AttFuse xuOPV2VOpenBenchmark2022
51.92 36.89 22.96 37.27 40.20 11.54 0.04 18.58
DiscoNet liLearningDistilledCollaboration2021
39.83 53.19 37.37 43.46 33.74 33.68 23.61 30.34
V2VNet wangV2VNetVehicletoVehicleCommunication2020
41.33 51.32 37.21 43.29 35.05 36.06 22.10 31.07
CoBEVT xuCoBEVTCooperativeBird2022
52.16 46.58 38.07 45.60 25.43 22.69 7.16 18.42
Ours
55.05 50.84 39.48 46.96 42.01 41.20 28.33 37.18
Domain Type
Rain Night
AP@IoU(%)
Vehicle Road Lane Average Vehicle Road Lane Average
AttFuse xuOPV2VOpenBenchmark2022
47.93 29.15 0.14 25.74 22.58 9.68 2.22 11.49
DiscoNet liLearningDistilledCollaboration2021
36.83 45.48 29.96 37.42 22.62 5.41 9.42 12.48
V2VNet wangV2VNetVehicletoVehicleCommunication2020
41.38 43.47 29.72 38.19 22.43 14.52 4.73 13.89
CoBEVT xuCoBEVTCooperativeBird2022
50.67 29.33 28.24 36.08 5.53 20.16 2.15 9.28
Ours
49.77 36.52 30.80 39.03 25.78 21.61 12.73 20.04
@@ IV-B Implementation Details
Our model is built on PyTorch and trained on two RTX4090 GPUs utilizing the AdamW optimizer. In the training process, the initial learning rate is set to $2\times 10^{-4}$ and decays by an exponential factor of $1\times 10^{-2}$ with a cosine annealing learning rate scheduler. During meta-consistency training, the meta-step size is set to $1\times 10^{-3}$. The hyperparameter $\alpha$ is empirically set to 0.01 to avoid artifacts on the transformed images. The balance weight $\beta$ of consistency loss is set to 0.1. We set the communication range of CAVs to 70m, and all the CAVs outside the communication range from the ego CAV are ignored. In addition, we first scale the camera images into $512\times 512$ pixels, then we employ CVT zhouCrossviewTransformersRealtime2022 as our backbone to extract the features of each CAV, we employ FuseBEVT xuCoBEVTCooperativeBird2022 to fuse the features from different CAVs, and after that we employ a 6 layer convolutional neural network as the decoder to generate the final BEV segmentation map.
@@ IV-C Comparison Analysis
Experimental setting. In our experiment, we first train the all the models in the original OPV2V dataset for 90 epochs with batch size equal to 1, then we continue to train the model on the synthesized datasets for 20 epochs. During continuing training on the synthesized datasets, we follow the convention in domain generalization literature by employing the leave-one-domain-out strategy. Specifically, we train on $K-1$ distributed source domains and evaluate on the remaining unseen target domain. This approach yields four distinct generalization settings for the collaborative BEV segmentation task.
In assessing the performance of our BEV segmentation rendered by our proposed approach, we compare it with several methods as follows: 1) AttFuse xuOPV2VOpenBenchmark2022 : employs self-attention to fuse the features derived from collaborative vehicles; 2) V2VNet wangV2VNetVehicletoVehicleCommunication2020 : leverages a spatially cognizant graph neural network to consolidate information garnered from all proximate CAVs; 3) DiscoNet liLearningDistilledCollaboration2021 : deploys a teacher-student framework to instruct the model via knowledge distillation; 4) CoBEVT xuCoBEVTCooperativeBird2022 : fuses camera features from a myriad of views and agent data within an underlying transformer framework. For the implementation of these methods, we follow the original papers and use the open code provided by the authors and establish them in the collaborative BEV segmentation setting with the same decoder.
@@
Comparison results. The results of our experiments are summarized in Table I, showcasing the domain generalization performance across different environmental conditions. It is evident from the table that our proposed method outperforms the state-of-the-art techniques across all tested domains, affirming the robustness and efficacy of our approach. In the Sunny domain, our method achieves the highest average precision (AP) of 46.96%, a substantial improvement over the next best performance of 45.60% by CoBEVT. Similarly, in the Fog domain, our method again tops the list with an AP of 37.18%, demonstrating a notable gain over the 31.07% achieved by V2VNet. Moreover, in challenging Rain and Night domains, our approach continues to exhibit superior performance. In the Rain domain, we achieve an AP of 39.03%, outclassing the 38.19% by V2VNet, and in the Night domain, our method significantly surpasses others with an AP of 20.04%, with the next best being 13.89% by V2VNet. The comparative analysis elucidates the benefit of employing a unified domain generalization framework for training and inference in collaborative perception. Our method’s ability to generalize across various domains validates the efficacy of the employed amplitude augmentation and meta-consistency training scheme. The results also underline the importance of intra-system domain alignment during the inference phase, significantly reducing the domain discrepancy among CAVs and consequently enhancing the collaborative perception performance across different environmental conditions.
@@
Figure 8: Visualization of the BEV segmentation results of different methods. In the first row, The four images are from ego vehicle’s cameras that in four different angles. In the second row, the five BEV maps are ground truth and generated from different method: column (a) is the ground truth, (b) is generated from AttFuse method, (c) is from DiscoNet, (d) is from V2VNet, (e) is predicted by our method. The third row and forth row are another scene. Compared with other methods, our method shows robust performance in different scenes.
@@
Qualitative analysis. In order to provide a qualitative comparison across different methods, we visualize some BEV segmentation results of different methods (including AttFuse, DiscoNet, V2VNet) in Fig. 8. As we can see in the BEV prediction maps, our model yields perception results that stand out in terms of both comprehensiveness and accuracy when compared with other methods. In both scenes, the AttFuse method exhibits significant omissions in road surface and lane. DiscoNet and V2VNet also more or less fail to segment the complete road and lane. As for the vehicle class, these three compared methods occasionally miss segments and display ambiguous boundaries. In contrast, as seen in Fig. 8 (e), our method can segment the road, lane and vehicle class more accurately and comprehensively. These results can show the superiority of our method in the collaborative perception task.
@@ IV-D Ablation Study
Figure 9: Effect of the meta-consistency training ”w/” means with meta-consistency training, ”w/o” means without meta-consistency training.
Figure 10: Visualization of the consistency loss $\mathcal{L}_{cons}$, meta-test loss and meta-train loss.
Effect of amplitude augmentation. In order to evaluate the effectiveness of the proposed amplitude augmentation, we set the sunny dataset as the unseen domain, and test its performance. Then we conduct an experiment depicted in Table II. We observe that by simply removing the amplitude augmentation, the performance degrade. The IoU accuracy of vehicles, road, lane are decreased by 9.51%, 5.58%, 3.51%, respectively. To further indicate the effect of the AmpAug mechanism and show that AmpAug can increase the diversity of the source domain. We use t-SNE111t-SNE: t-distributed stochastic neighbor embedding is a statistical method for visualizing high-dimensional data by giving each datapoint a location in a two or three-dimensional map. van2008visualizing to visualize the distribution of the original images and the generated images in Fig. 11. Specifically, the pink points denote the local data, the other blue points denote the transformed data that are generated with amplitude spectrum from our constructed dataset $\mathcal{D}_{t}$ proposed in Sec. III-B. From the observation, we can see that the distribution of the transformed images is more diverse than the original images. This indicates that the AmpAug mechanism can increase the diversity of the source domain by augmenting the low-level distribution of the images, thereby making the model to see more low-level variations. This can prevent the model from overfitting to the source domain and help the model to learn the domain-invariant features.
Figure 11: Visualization of the t-SNE embedding for the original dataset (pink points) and the corresponding transformed images by amplitude augmentation dataset (blue points).
TABLE II: Ablation Study Results of the Amplitude Augmentation. “w/” means the with amplitude augmentation, “w/o” means the without amplitude augmentation.
AP@IoU(%) Vehicle Road Lane Average
@@
TABLE III: Ablation Study Results of the Intra-System Domain Alignment. “w/” means the with intra-system domain alignment, “w/o” means the without intra-system domain alignment.
AP@IoU (%) Vehicles Road Lane Average
@@
V2VNet wangV2VNetVehicletoVehicleCommunication2020 w/o 41.33 51.32 37.21 43.29
@@
Effect of meta-consistency training. In order to evaluate the effectiveness of the proposed meta-consistency training, we conduct an experiment depicted in Fig. 9. We observe that the meta-consistency training does improve the performance of the model: the IoU accuracy of vehicles, road, lane are improved by 2.56%, 1.08% and 0.26%, respectively. Without our method, the model could not learn the domain-invariant features, which leads to the performance degradation. Notably, the performance of vehicle class is more sensitive to the domain discrepancy than the other classes, because vehicles are more dynamic than the other classes. The results demonstrate that our meta-consistency training can improve the performance of collaborative perception and make the model more robust to the domain shift.
@@
To further analyze the effectiveness of the meta-consistency training, we visualize the consistency loss, denoted as $\mathcal{L}_{cons}$ in Fig. 10. It can be observed that despite the consistency loss $\mathcal{L}_{cons}$ has some fluctuations, it shows the decreasing trend during the training process. This suggests that the consistency loss can make the latent representations more similar, thereby aiding the model to learn the domain-invariant features. In addition, the meta-test loss and meta-train loss also show a decreasing trend. The meta-test loss has a large fluctuation at the beginning during the training, but it gradually levels off around 0.06.
@@
Effect of intra-system domain alignment. In order to evaluate the effect and the university of our intra-system domain alignment mechanism, we conduct an experiment with different collaborative perception framework, including AttFuse xuOPV2VOpenBenchmark2022 , V2VNet wangV2VNetVehicletoVehicleCommunication2020 , and DiscoNet liLearningDistilledCollaboration2021 . Before evaluating on these different collaborative perception frameworks, we first implement our method on each framework. The results are shown in Table III. We can see that our intra-system domain alignment mechanism can generally improve the performance of the baselines, especially for the vehicle class. Specifically, for the AttFuse method, the IoU accuracy of vehicles, road and lane are improved by 1.24%, 4.23%, 2.64%, respectively. For the V2VNet, the IoU accuracy of vehicles, road and lane are improved by 3.52%, 2.17%, 1.92%, respectively. In addition, as for our method, we simply remove this component and evaluate, we can see that the performance of the model will be degraded. These results can demonstrate the effect of our intra-system domain alignment mechanism that can generally improve the performance of the collaborative perception framework by reducing the distribution heterogeneity among the data in different vehicles.
V Conclusion

### v3

@@ IV-A Datasets and Metrics
To assess the domain generalization performance of our proposed method, it is essential to evaluate it across various datasets exhibiting domain discrepancies. However, in the domain of collaborative BEV segmentation, there are no public available datasets suitable for this purpose. Consequently, we have taken the initiative to synthesize four distinct datasets, each characterized with unique domain discrepancies, drawing from the OPV2V dataset xuOPV2VOpenBenchmark2022 as the foundation. These datasets include variations, including foggy weather, rainy weather, night conditions, and bright sunny scenarios. The specific attributes and details of these synthesized datasets are elaborated below.
OPV2V. In our study, we leverage the OPV2V dataset xuOPV2VOpenBenchmark2022 , a comprehensive dataset curated for integrated perception using V2V communications. This dataset, amassed through the CARLA simulator dosovitskiyCARLAOpenUrban2017 and OpenCDA xuOpenCDAOpenCooperative2021 , comprises 11,464 frames of LiDAR point clouds and images, with each frame featuring the minimum of 2 and the maximum of 7 interconnected vehicles. It encompasses 73 distinct scenarios, each averaging a duration of 25 seconds.
Sunny. The initial task in our setting is driving on a sunny day. As the OPV2V dataset is collected in the daytime, so we sample a portion of OPV2V as our sunny dataset.
Fog. The second task is driving on a foggy day. We synthesize a foggy-weather dataset with different densities of fog (e.g., dense fog and light fog) as shown in Fig. 5 and 6, respectively. The foggy weather is generated from the atmospheric scattering model narasimhanVisionAtmosphere2008 .
Rain. In addition, we synthesize a rainy-weather dataset with different densities of rain (e.g., heavy rain and light rain) as shown in Fig. 7. Specifically, the dataset is generated by utilizing RainMix, which is a novel data augmentation method. In order to generate the rainy dataset, we first sample a rain map from a public dataset of real rain streaks gargPhotorealisticRenderingRain2006 . We then apply random transformations to the rain map using the RainMix technique. These transformations include rotation, zooming, translation, and shearing, which are randomly sampled and combined. Lastly, the rain maps after transformation are merged with the original source domain images.
Night. Finally, as shown in Fig. 8, we synthesize a night dataset by leveraging a day-to-night generative adversarial network (GAN) zhuUnpairedImageImageTranslation2017 , which can render the daytime images to night images. The four datasets under different weather conditions contain 13,920 images in total, and each dataset contains 3,480 images.
@@
Domain Type Sunny Fog
AP@IoU(%) Vehicle Road Lane Average Vehicle Road Lane Average
AttFuse xuOPV2VOpenBenchmark2022 51.92 36.89 22.96 37.27 40.20 11.54 0.04 18.58
V2VAM liLearningVehicletoVehicleCooperative2023 48.85 35.14 20.08 34.69 41.83 21.56 8.04 23.81
F-Cooper chenFCooperFeatureBased2019 42.27 32.34 25.01 33.17 32.45 12.83 9.04 18.10
Where2Comm huWhere2commCommunicationefficientCollaborative2024 44.13 30.50 24.89 33.17 26.70 5.91 9.00 13.87
V2X-ViT xuV2XViTVehicletoEverythingCooperative2022a 37.86 36.35 24.91 33.04 25.79 23.13 10.19 19.70
DiscoNet liLearningDistilledCollaboration2021 39.83 53.19 37.37 43.46 33.74 33.68 23.61 30.34
V2VNet wangV2VNetVehicletoVehicleCommunication2020a 41.33 51.32 37.21 43.29 35.05 36.06 22.10 31.07
CoBEVT xuCoBEVTCooperativeBird2023 52.16 46.58 38.07 45.60 25.43 22.69 7.16 18.42
Ours 55.05 50.84 39.48 48.46 42.01 41.20 28.33 37.18
Domain Type Rain Night
AP@IoU(%) Vehicle Road Lane Average Vehicle Road Lane Average
AttFuse xuOPV2VOpenBenchmark2022 47.93 29.15 0.14 25.74 22.58 9.68 2.22 11.49
V2VAM liLearningVehicletoVehicleCooperative2023 46.62 26.98 18.90 30.83 13.11 11.54 3.87 9.51
F-Cooper chenFCooperFeatureBased2019 42.50 26.32 18.52 29.11 14.59 9.21 5.27 9.66
Where2Comm huWhere2commCommunicationefficientCollaborative2024 40.26 28.31 24.26 30.94 11.78 11.91 7.99 10.56
V2X-ViTxuV2XViTVehicletoEverythingCooperative2022a 38.10 26.05 16.51 26.89 15.21 11.27 6.97 11.15
DiscoNet liLearningDistilledCollaboration2021 36.83 45.48 29.96 37.42 22.62 5.41 9.42 12.48
V2VNet wangV2VNetVehicletoVehicleCommunication2020a 41.38 43.47 29.72 38.19 22.43 14.52 4.73 13.89
CoBEVT xuCoBEVTCooperativeBird2023 50.67 29.33 28.24 36.08 5.53 20.16 2.15 9.28
Ours 49.77 36.52 30.80 39.03 25.78 21.61 12.73 20.04
@@ IV-B Implementation Details
Our model is built on PyTorch and trained on two RTX4090 GPUs utilizing the AdamW optimizer. In training process, the initial learning rate is set to $2\times 10^{-4}$ and decays by an exponential factor of $1\times 10^{-2}$ with a cosine annealing learning rate scheduler. During meta-consistency training, the meta-step size is set to $1\times 10^{-3}$. The hyperparameter $\alpha$ is empirically set to 0.01 to avoid artifacts on the transformed images. The balance weight $\beta$ of consistency loss is set to 0.1. We set the communication range of CAVs to 70m, and all the CAVs outside the communication range from the ego CAV are ignored. In addition, we first scale the camera images into $512\times 512$ pixels, then we employ CVT zhouCrossviewTransformersRealtime2022a as our backbone to extract the features of each CAV, we employ FuseBEVT xuCoBEVTCooperativeBird2023 to fuse the features from different CAVs, and after that we employ a six layer convolutional neural network (CNN) as the decoder to generate the final BEV segmentation map.
@@ IV-C Comparison Analysis
Experimental Setting. In our experiment, we first train all the models in the original OPV2V dataset for 90 epochs with batch size equal to 1, then we continue to train the model by our proposed meta-consistency training paradigm for 20 epochs. After that, we evaluate the performance on the synthesized datasets. In assessing the performance of our BEV segmentation rendered by our proposed approach, we compare it with several methods, including AttFuse xuOPV2VOpenBenchmark2022 , V2VAM liLearningVehicletoVehicleCooperative2023 , F-Cooper chenFCooperFeatureBased2019 , Where2Comm huWhere2commCommunicationefficientCollaborative2024 , V2X-ViT xuV2XViTVehicletoEverythingCooperative2022a , DiscoNet liLearningDistilledCollaboration2021 , V2VNet wangV2VNetVehicletoVehicleCommunication2020a , and CoBEVT xuCoBEVTCooperativeBird2023 . For the implementation of these methods, we follow the original papers and use the open code provided by the authors and establish them in the collaborative BEV segmentation setting with the same decoder.
@@
Comparison Results. The results of our experiments are summarized in Table I, showcasing the domain generalization performance across different environmental conditions. It is evident from the table that our proposed method significantly outperforms the state-of-the-art techniques across all test domains, affirming the robustness and efficacy of our approach. In the Sunny domain, our method achieves the highest average precision (AP) of 46.96%, a substantial improvement over the next best performance of 45.60% by CoBEVT. Similarly, in the Fog domain, our method again tops the list with an AP of 37.18%, demonstrating a notable gain over 31.07% achieved by V2VNet. Moreover, in challenging Rain and Night domains, our approach continues to exhibit superior performance. In the Rain domain, we achieve an AP of 39.03%, outclassing the 38.19% by V2VNet, and in the Night domain, our method significantly surpasses others with an AP of 20.04%, with the next best being 13.89% by V2VNet. The comparative analysis elucidates the benefit of employing a unified domain generalization framework for training and inference in CP. Our method’s ability to generalize across various domains validates the efficacy of the employed amplitude augmentation and meta-consistency training scheme. The results also underline the importance of the intra-system domain alignment during the inference phase, significantly reducing the domain discrepancy among CAVs and consequently enhancing the collaborative perception performance across different environmental conditions.
@@
Figure 9: Visualization of the BEV segmentation results of different methods. In the first row, The four images are from ego vehicle’s cameras from four different angles. In the second row, the five BEV maps are ground truth and generated from different method: column (a) is the ground truth, (b) is generated from AttFuse, (c) is from DiscoNet, (d) is from V2VNet, and (e) is predicted by our method. The third row and forth row are for another scene. Compared with other methods, our method shows robust performance in different scenes.
@@
Qualitative Analysis. In order to provide a qualitative comparison across different methods, we visualize some BEV segmentation results of different methods (including AttFuse, DiscoNet, V2VNet) in Fig. 9. As we observe in the BEV prediction maps, our model yields the perception results that stand out in terms of both comprehensiveness and accuracy when compared with other methods. In both scenes, AttFuse exhibits significant omissions in road surface and lane. DiscoNet and V2VNet more or less fail to segment the complete road and lane. As for the vehicle class, these three compared methods occasionally miss segments and display ambiguous boundaries. In contrast, as seen in Fig. 9 (e), our method can segment the road, lane, and vehicle class more accurately and comprehensively. These results can show the superiority of our method in the CP task.
@@ IV-D Ablation Study
Figure 10: Effect of the meta-consistency training. “w/” means with meta-consistency training, “w/o” means without meta-consistency training.
Figure 11: Visualization of the consistency loss $\mathcal{L}_{cons}$, meta-test loss, and meta-train loss. Subfigure (d) visualizes AP@IoU tested on the foggy and rainy domains with different epochs.
Effect of Amplitude Augmentation. In order to evaluate the effectiveness of the proposed amplitude augmentation, we set the sunny dataset to the unseen domain and test its performance. Then, we conduct an experiment depicted in Table II. We observe that by simply removing the amplitude augmentation, the performance degrades. The IoU accuracy of vehicles, road, and lane are decreased by 9.51%, 5.58%, and 3.51%, respectively. In addition, we conduct comparative experiments with conventional augmentation methods, RainMix gargPhotorealisticRenderingRain2006 and Color Augmentation (ColorAug) otaloraStainingInvariantFeatures2019 , to show the efficiency of AmpAug. In the experiment, we replace the AmpAug module with others, respectively. From Table II, we observe that our AmpAug can achieve better performance than the other two methods, further demonstrating the effectiveness of our AmpAug.
To further capture the effect of AmpAug and show that AmpAug can increase the diversity of the source domain. We use t-SNE111t-SNE: t-distributed stochastic neighbor embedding is a statistical method for visualizing high-dimensional data by giving each data point a location in a two or three-dimensional map. van2008visualizing to visualize the distribution of the original images and the generated images in Fig. 12. Specifically, the pink points denote the local data, the other blue points denote the transformed data that are generated with AmpAug from our constructed dataset $\mathcal{D}_{t}$ proposed in Sec. III-B. From the observation, we can see that the distribution of the transformed images is more diverse than the original images. This indicates that AmpAug can increase the diversity of the source domain by augmenting the low-level distribution of the images, thereby making the model see more low-level variations. This can prevent the model from overfitting the source domain and help the model learn the domain-invariant features.
Figure 12: Visualization of the t-SNE embedding for the original dataset (pink points) and the corresponding transformed images by AmpAug dataset (blue points).
TABLE II: Ablation Study and Comparative Results of AmpAug. “w/” means with AmpAug, “w/o” means without AmpAug.
Methods Vehicles Road Lane Average
@@
TABLE III: Ablation Study Results of the Intra-System Domain Alignment. “w/” means with intra-system domain alignment, “w/o” means without intra-system domain alignment.
Methods Vehicles Road Lane Average
@@
V2VNet wangV2VNetVehicletoVehicleCommunication2020a w/o 41.33 51.32 37.21 43.29
@@
Effect of Meta-consistency Training. In order to evaluate the effectiveness of the proposed meta-consistency training, we conduct an experiment depicted in Fig. 10. We observe that the meta-consistency training does improve the performance of the model: the IoU accuracy of vehicles, road, and lane are improved by 2.56%, 1.08% and 0.26%, respectively. Without our method, the model could not learn the domain-invariant features, which leads to performance degradation. Notably, the performance of the vehicle class is more sensitive to the domain discrepancy than the other classes, because vehicles are more dynamic than the other classes. The results demonstrate that our meta-consistency training can improve the performance of CP and make the model more robust to the domain shift.
@@
To further analyze the effectiveness of the meta-consistency training, we visualize the consistency loss, denoted as $\mathcal{L}_{cons}$ in Fig. 11. It can be observed that despite the consistency loss $\mathcal{L}_{cons}$ has some fluctuations, it shows a decreasing trend during the training process. This suggests that the consistency loss can make the latent representations more similar, thereby aiding the model to learn the domain-invariant features. In addition, the meta-testing loss and meta-training loss also show a decreasing trend. The meta-testing loss has a large fluctuation at the beginning of the training, but it gradually levels off around 0.06. In addition, we also evaluate mIoU in the test set in fog and rain with different epochs, as shown in Fig. 11(d), we observe that the AP levels off when the number of epochs is around 100, which reflects the convergence of the loss.
@@
Effect of Intra-system Domain Alignment. In order to evaluate the effect and the university of our intra-system domain alignment mechanism, we conduct an experiment with different collaborative perception frameworks, including AttFuse xuOPV2VOpenBenchmark2022 , V2VNet wangV2VNetVehicletoVehicleCommunication2020a , and DiscoNet liLearningDistilledCollaboration2021 . Before evaluating these different collaborative perception frameworks, we first implement our method on each framework. The results are shown in Table III. We observe that our intra-system domain alignment mechanism can generally improve the performance of the baselines, especially for the vehicle class. Specifically, for AttFuse, the IoU accuracy of vehicles, road, and lane are improved by 1.24%, 4.23%, and 2.64%, respectively. For V2VNet, the IoU accuracy of vehicles, road, and lane are improved by 3.52%, 2.17%, and 1.92%, respectively. In addition, as for our method, we simply remove this component and evaluate it, we can see that the performance of the model will be degraded. These results can demonstrate the effect of our intra-system domain alignment mechanism that can generally improve the performance of our CP framework by reducing the distribution heterogeneity among the data in different vehicles.
Collective Ablation Study. In this section, we further analyze the collective impact of these techniques on the performance of our overall framework and provide a deeper understanding of their respective contributions and synergistic effects within the CP framework. As shown in Tab. IV, we conduct an ablation study of our proposed method. We compare the performance of our method with the vanilla method, which removes the three techniques in our method, including AmpAug, meta-consistency training, and intra-system domain alignment. The results show that our method outperforms the vanilla method in most cases, which indicates that each component contributes to the overall performance.
TABLE IV: Collective Ablation Study Results of the Whole Techniques.. ‘Vanilla’ represents that we remove the three techniques in our method, including AmpAug, meta-consistency training, and intra-system domain alignment.
Method Domain Vehicles Road Lane Average
Vanilla Sunny 46.67 47.33 36.66 43.55
Fog 44.23 35.37 26.88 35.56
Rain 42.35 33.10 29.34 34.93
Night 20.15 15.61 9.57 15.11
Ours Sunny 55.05 50.84 39.48 46.96
Fog 42.01 41.20 28.33 37.18
Rain 49.77 36.52 30.80 39.03
Night 25.78 21.61 12.73 20.04
V Limitations and Failure Cases
Our work introduces a novel framework to tackle the full-scene domain generalization problem in multi-agent collaborative BEV segmentation. While domain generalization is inherently challenging, and no existing works, including ours, can completely solve it for all scenarios, and thus, we acknowledge certain limitations and failure cases that require further investigation. For instance, although our method significantly mitigates domain shift, it struggles when the shift is substantial or when the target domain starkly differs from the source domain. A specific example is the performance discrepancy between day and night scenarios. Despite the effectiveness of our techniques, the model’s performance at night remains inferior to daytime results. We are committed to refining our approach and addressing these challenges in future work.
VI Conclusion